The sum of two numbers is 32. One of the numbers is 4 less than 5 times the other. Find the two numbers?
Answer by JulietG(1812)

A + B = 32
A = 5B -4
Substitute the known value of A from the second equation into the first.
5B-4 + B = 32
6B - 4 = 32
Add 4 to each side
6B = 36
Divide each side by 6
B = 6
.
If B is 6, then A is 32-6, or 26.